# DressCode: Autoregressively Sewing and Generating Garments from Text Guidance

### v1

@@
Abstract
Apparel’s significant role in human appearance underscores the importance of garment digitalization for digital human creation. Recent advances in 3D content creation are pivotal for digital human creation. Nonetheless, garment generation from text guidance is still nascent. We introduce a text-driven 3D garment generation framework, DressCode, which aims to democratize design for novices and offer immense potential in fashion design, virtual try-on, and digital human creation. For our framework, we first introduce SewingGPT, a GPT-based architecture integrating cross-attention with text-conditioned embedding to generate sewing patterns with text guidance. We also tailored a pre-trained Stable Diffusion for high-quality, tile-based PBR texture generation. By leveraging a large language model, our framework generates CG-friendly garments through natural language interaction. Our method also facilitates pattern completion and texture editing, simplifying the process for designers by user-friendly interaction. With comprehensive evaluations and comparisons with other state-of-the-art methods, our method showcases the best quality and alignment with input prompts. User studies further validate our high-quality rendering results, highlighting its practical utility and potential in production settings.
1 Introduction
@@
Recent years have witnessed tremendous progress in text-driven asset generation, triggered by the large-scale language models [1, 42]. It democratizes the accessible text-driven creation of diverse assets for novices, including images [45], generic 3D objects [28, 38], human hair [70], face or body [24, 67]. What is still missing is the garment. Moreover, naively applying avatar or general generation [24, 38] for the garment category is suboptimal since they turn to generate mesh or neural fields that are incompatible with digital garment production workflow.
@@
In contrast, for our graphics community, the dominant representation of garments is sewing patterns, which facilitates both physical simulation and animation in a CG-friendly fashion [8, 2]. For sewing pattern generation, early methods [58, 6] only simply partial modules in the workflow like parsing or draping to 3D. With the consolidation of more advanced datasets [21], very recent work enables sewing pattern generation from point clouds [22] or images [26]. However, they largely overlook the generation through more natural language interactions, let alone handling the vivid generation with desired texture patterns or physically-based materials.
@@
In this paper, we propose DressCode, a 3D garment generation framework that generates high-quality garments via natural language interaction. As illustrated in Figure LABEL:fig:teaser, DressCode allows users to customize garments with preferred sewing patterns and physically-based texture details through text interaction. The resulting garments can be seamlessly integrated with CG pipelines, supporting post-editing and animation while ensuring high-quality rendering.
@@
Notably, the garments’ highly symmetric and structured nature, with uniform panels and stitching, leads to convenient conversion from sewing patterns to discrete “codes”. To this end, inspired by powerful language generation apt for this nature, we introduce SewingGPT, a GPT-based architecture for sewing pattern generation. Specifically, we adopt a novel quantization process to translate the sewing patterns into token sequences and subsequently utilize a decoder-only Transformer with text-conditioned embedding for token prediction. For effective training, we apply GPT-4V [1] on the existing dataset [21] to detail diverse garment types and shapes with rich text prompts. Once trained, our SewingGPT autoregressively generates quantized sewing patterns with efficient text interactions, which have been unseen before.
@@
To achieve high-quality garment rendering, we progressively tailor a pre-trained Stable Diffusion [45] to generate tile-based physically-based rendering (PBR) textures from text prompts. We first fine-tune the U-Net architeture [46] of the diffusion model within latent space to generate the diffuse attribute, and then fine-tune the various VAE [20] decoders to generate normal and roughness maps separately.
@@
2 Related Work
Figure 1: Overview of our SewingGPT pipeline. We quantize sewing patterns to the sequence of tokens and adopt a GPT-based architecture to generate the tokens autoregressively. Our SewingGPT enables users to generate highly diverse and high-quality sewing patterns under text prompt guidance.
@@ 2 Related Work
Sewing pattern representation is vital for garment modeling. Recent studies have delved into sewing pattern reconstruction [37, 19, 64, 60, 48, 55], generation [50], draping [13, 23, 6] and editing [5, 58, 39]. Early research [12] employs a search in a pre-defined database of 3D garment parts for garment reconstruction. Studies [19, 54, 64] use parametric sewing patterns, optimizing them for garment reconstruction from images. [60] advances this by applying deep learning to discern a shape space for sewing patterns and various input modalities. [4, 48] utilize surface flattening for sewing pattern reconstruction from 3D human models.
Recently, some work [71, 15, 22] adopt data-driven approaches for reconstruction. [15] utilizes a deep network with surface flattening for sewing pattern reconstruction from 3D geometries. [21] generates a sewing pattern dataset covering a wide range of garment shapes and topologies. NeuralTailor [22] offers advanced sewing pattern reconstruction using a hybrid network to predict garment panels and stitching information from point cloud input. [11] introduces a CNN-based model capable of predicting garment panels from single images, using PCA to simplify the panel data structure. [26] creates a comprehensive dataset featuring diverse garment styles and human poses and introduces a two-level Transformer network, achieving state-of-the-art sewing pattern reconstruction from single images. While these solutions yield notable results, a gap persists in user-friendliness and practicality compared to direct communication of outcomes through natural language. Furthermore, prior studies have largely overlooked generating garment color, texture, and material, essential elements for creating high-quality garments.
@@
Recent breakthroughs in the text-to-image domain [16, 68, 45] have enhanced interest in text-guided 3D content generation. Early work [18] introduces a text-to-3D method guided by CLIP [42]. [38, 59] present the Score Distillation Sampling (SDS) algorithm, elevating pre-trained 2D diffusion models for the 3D generation. [33] optimizes Neural Radiance Fields (NeRF) [34] in the latent space. [25, 10] optimize efficient mesh representations [35, 49] for higher quality generation. [47] integrates 3D awareness into 2D diffusion for improving text-to-3D generation consistency. Subsequent studies [9, 31, 44] focus on texturing pre-existing meshes, balancing speed and quality. Despite their innovations, SDS-based methods faced challenges with over-saturation. [57] proposes a novel method to refine mesh textures for more realistic generations. ProlificDreamer [61] introduces the Variational Score Distillation (VSD) method to mitigate over-saturation effectively. Furthermore, several studies [52, 65, 69, 29] explored 3D generation using multi-view diffusion. Concurrently, some research [17, 27, 28, 30, 40, 43, 51, 56, 63, 62, 32] concentrates on reconstructing 3D content from a single image through distillation, achieving high-fidelity textured meshes from 2D diffusion priors. Additionally, some studies [66, 14, 36, 53] delve into shape generation. [66] generates high-quality 3D shapes with Unsigned Distance Field (UDF) through Diffusion models. [36, 53] employ autoregressive models for mesh structure generation. Although some general object generation methods [41, 34, 38, 61, 59, 66] can produce garments, their practicality in CG environments is limited. These 3D outputs, mostly mesh-based or derived from implicit fields, lack adaptability for fitting different bodies and layering multiple garments, common needs in garment design. Furthermore, the textures, typically produced via optimization or multi-view reconstruction, are often low-resolution and blurry, neglecting the structured UV mapping of garments, resulting in poor topology challenging for subsequent CG processing.
3 Sewing Pattern Generation
Figure 2: Overview of our entire DressCode pipeline for customized garment generation. We employ a large language model to obtain shape prompts and texture prompts with natural language interaction and utilize the SewingGPT and a fine-tuned Stable Diffusion for high-quality and CG-friendly garment generation.
@@ 3 Sewing Pattern Generation
3.1 Sewing Pattern Quantization
@@
We utilize the sewing pattern templates from [22], which cover a wide variety of garment shapes. Each sewing pattern includes $N_{P}$ panels $\{P_{i}\}_{i=1}^{N_{P}}$ and stitching information $S$. Each panel $P_{i}$ forms a closed 2D polygon with $N_{i}$ edges. Each edge $E$ consists of two 2D points $(v_{x},v_{y})$, the edge’s start, and $(c_{x},c_{y})$, the control point of the Bezier curve. The 3D placement of each panel is indicated by rotation quaternion $(q_{1},q_{2},q_{3},q_{4})$ and translation vector $(t_{1},t_{2},t_{3})$. For stitching information, each edge has a feature vector $(u_{0/1},s_{1},s_{2},s_{3})$ where $u_{0/1}$ is a binary flag indicating whether there is a stitch on this edge, and $(s_{1},s_{2},s_{3})$ is a stitch tag based on 3D placement of the edge.
@@
For each panel, we first standardize and quantize all parameters, subsequently converting them into tokens. These tokens are then formed into a linear sequence. We set a maximum limit, denoted as $K$, for the number of edges in each panel. To maintain a uniform token count across panels, we apply zero-padding to panels with $N_{i}<K$. Subsequently, all panels are flattened and merged into a single sequence, starting with a start token and ending with an end token. Owing to the uniformity of token counts for each panel, inserting padding tokens between panels is not required. Consequently, the resultant sequence, denoted as $\mathcal{F^{\text{seq}}}$, spans a length of $L_{t}=(8K+7)N_{P}$, with each token denoted by $f_{n}$ for $n=1,\ldots,L_{t}$. Eventually, we can represent it as
$\mathcal{F^{\text{seq}}}=\{\mathcal{T}(P_{i},S)\}_{i=1}^{N_{P}},$ (1)
where $\mathcal{T}$ denotes the quantization function.
3.2 Generation with Autoregressive Model
Utilizing GPT-based architectures, we adopt a decoder-only transformer to generate token sequences for sewing patterns. Inspired by Polygen [36], we design the triple embedding for each input token: positional embedding, denoting which panel it belongs; parameter embedding, classifying the token as edge coordinates, rotation, translation, or stitching feature vectors; and value embedding for the quantized sewing pattern values. The source tokens are then input into the transformer decoder to predict the probability distribution of the next potential token at each step. Consequently, the objective is to maximize the log-likelihood of the training sequences:
$\mathcal{L}=-\prod_{i=0}^{L_{t}}p\left(f_{i}\mid f_{<i};\theta\right).$ (2)
@@ 3.2 Generation with Autoregressive Model
To guide the sewing pattern generation, our model integrates cross-attention with text-conditioned embeddings h. Initially, we utilize the CLIP model to obtain the CLIP embedding, which is then projected into a feature embedding through a compact Multilayer Perceptron (MLP). Subsequently, the Transformer decoder conducts cross-attention with the feature embedding, facilitating condition-specific token generation.
Figure 3: Examples of our data captions. We utilize the rendered images and ask GPT-4V with the designed prompt for detailed captions.
3.3 Implementation Details

### v2

@@
Abstract.
Apparel’s significant role in human appearance underscores the importance of garment digitalization for digital human creation. Recent advances in 3D content creation are pivotal for digital human creation. Nonetheless, garment generation from text guidance is still nascent. We introduce a text-driven 3D garment generation framework, DressCode, which aims to democratize design for novices and offer immense potential in fashion design, virtual try-on, and digital human creation. For our framework, we first introduce SewingGPT, a GPT-based architecture integrating cross-attention with text-conditioned embedding to generate sewing patterns with text guidance. We also tailored a pre-trained Stable Diffusion for high-quality, tile-based PBR texture generation. By leveraging a large language model, our framework generates CG-friendly garments through natural language interaction. Our method also facilitates pattern completion and texture editing, streamlining the design process through user-friendly interaction. This framework fosters innovation by allowing creators to freely experiment with designs and incorporate unique elements into their work, thereby igniting new ideas and artistic possibilities. With comprehensive evaluations and comparisons with other state-of-the-art methods, our method showcases the best quality and alignment with input prompts. User studies further validate our high-quality rendering results, highlighting its practical utility and potential in production settings. Our project page is https://IHe-KaiI.github.io/DressCode/.
Garment Generation, Sewing Patterns, Autoregressive Model
copyright: rightsretainedjournal: TOGjournalyear: 2024journalvolume: 43journalnumber: 4publicationmonth: 7doi: 10.1145/3658147
Figure 1. Our DressCode generates CG-friendly customized garments with sewing patterns and PBR textures under natural text guidance, enabling post-editing, animation, and high-quality rendering.
1. Introduction
@@
Recent years have witnessed tremendous progress in text-driven asset generation, triggered by the large-scale language models (Achiam et al., 2023; Radford et al., 2021). It democratizes the accessible text-driven creation of diverse assets for novices, including images (Rombach et al., 2022), generic 3D objects (Liu et al., 2023b; Poole et al., 2022), human hair (Zhou et al., 2023), face or body (Liao et al., 2023; Zhang et al., 2023a). What is still missing is the garment. Moreover, naively applying avatar or general generation (Liao et al., 2023; Poole et al., 2022) for the garment category is suboptimal since they turn to generate mesh or neural fields that are incompatible with digital garment production workflow.
@@
In contrast, for our graphics community, the dominant representation of garments is sewing patterns, which facilitates both physical simulation and animation in a CG-friendly fashion (Blender Foundation, 2022; Autodesk, INC., 2019). For sewing pattern generation, early methods (Umetani et al., 2011; Berthouzoz et al., 2013) only use simple partial modules in the workflow like parsing or draping to 3D. With the consolidation of more advanced datasets (Korosteleva and Lee, 2021), very recent work enables sewing pattern generation from point clouds (Korosteleva and Lee, 2022) or images (Liu et al., 2023d). (Korosteleva and Sorkine-Hornung, 2023) design a DSL for sewing pattern construction, allowing parameterization and algorithmic support. However, they largely overlook the generation through more natural language interactions, let alone handling the vivid generation with desired texture patterns or physically-based materials, which could significantly speed up the preliminary stage of garment design. Furthermore, novices without professional skills in design or complex design software need to describe their ideas only in text prompts to transform these ideas into designs. This significantly lowers the design barriers, allowing more people to participate in the creative process. Most importantly, generative models introduce diversity to designed text prompts, generate varied types of garments conforming to the prompts, and stimulate designer creativity.
@@
In this paper, we propose DressCode, a 3D garment generation framework that generates high-quality garments via natural language interaction. As illustrated in Figure 1, DressCode allows users to customize garments with preferred sewing patterns and physically-based texture details through text interaction. The resulting garments can be seamlessly integrated with CG pipelines, supporting post-editing and animation while ensuring high-quality rendering.
@@
Notably, the garments’ highly symmetric and structured nature, with uniform panels and stitching, leads to convenient conversion from sewing patterns to discrete “codes.” To this end, inspired by powerful language generation apt for this nature, we introduce SewingGPT, a GPT-based architecture for sewing pattern generation. Specifically, we adopt a novel quantization process to translate the sewing patterns into token sequences and subsequently utilize a decoder-only Transformer with text-conditioned embedding for token prediction. We utilize the pre-trained CLIP (Radford et al., 2021) model to encode prompts as conditional embeddings, benefiting from CLIP’s generalized capability in multimodal understanding. For effective training, we apply GPT-4V (Achiam et al., 2023) on the existing dataset (Korosteleva and Lee, 2021) to detail diverse garment types and shapes with rich text prompts. Once trained, our SewingGPT autoregressively generates quantized sewing patterns with efficient text interactions, which have been unseen before.
@@
To achieve high-quality garment rendering, we progressively tailor a pre-trained Stable Diffusion model (Rombach et al., 2022) to generate tile-based physically-based rendering (PBR) textures from text prompts. We first fine-tune the U-Net architecture (Ronneberger et al., 2015) of the diffusion model within latent space to generate the diffuse attribute, and then fine-tune the various VAE (Kingma and Welling, 2013) decoders to generate normal and roughness maps separately.
@@
2. Related Work
@@
Sewing pattern representation is vital for garment modeling. Recent studies have delved into sewing pattern reconstruction (Pietroni et al., 2022; Jeong et al., 2015; Yang et al., 2018; Wang et al., 2018; Sharp and Crane, 2018; Su et al., 2022), generation (Shen et al., 2020), draping (De Luigi et al., 2023; Li et al., 2023b; Berthouzoz et al., 2013) and editing (Bartle et al., 2016; Umetani et al., 2011; Qi et al., 2023). Early research (Chen et al., 2015) employs a search in a pre-defined database of 3D garment parts for garment reconstruction. Studies (Jeong et al., 2015; Su et al., 2020; Yang et al., 2018) use parametric sewing patterns, optimizing them for garment reconstruction from images. (Wang et al., 2018) advances this by applying deep learning to discern a shape space for sewing patterns and various input modalities. (Bang et al., 2021; Sharp and Crane, 2018) utilize surface flattening for sewing pattern reconstruction from 3D human models.
Recently, some work (Zhu et al., 2020; Goto and Umetani, 2021; Korosteleva and Lee, 2022) adopt data-driven approaches for reconstruction. (Goto and Umetani, 2021) utilizes a deep network with surface flattening for sewing pattern reconstruction from 3D geometries. (Korosteleva and Lee, 2021) generates a sewing pattern dataset covering a wide range of garment shapes and topologies. NeuralTailor (Korosteleva and Lee, 2022) offers advanced sewing pattern reconstruction using a hybrid network to predict garment panels and stitching information from point cloud input. (Chen et al., 2022) introduces a CNN-based model capable of predicting garment panels from single images, using PCA to simplify the panel data structure. (Liu et al., 2023d) creates a comprehensive dataset featuring diverse garment styles and human poses and introduces a two-level Transformer network, achieving state-of-the-art sewing pattern reconstruction from single images. (Korosteleva and Sorkine-Hornung, 2023) designs the first DSL for garment modeling, enabling users to do rich garment designs using interchangeable, parameterized components. (Li et al., 2023a) proposes a novel approach to recover garment materials and patterns with optimization using differentiable simulation. While these solutions yield notable results, a gap persists in user-friendliness and practicality compared to direct communication of outcomes through natural language. Furthermore, prior studies have largely overlooked generating garment color, texture, and material, essential elements for creating high-quality garments.
@@
Recent breakthroughs in the text-to-image domain (Ho et al., 2020; Zhang et al., 2023b; Rombach et al., 2022) have enhanced interest in text-guided 3D content generation. Early work (Jain et al., 2022) introduces a text-to-3D method guided by CLIP (Radford et al., 2021). (Poole et al., 2022; Wang et al., 2023a) present the Score Distillation Sampling (SDS) algorithm, elevating pre-trained 2D diffusion models for the 3D generation. (Metzer et al., 2023) optimizes Neural Radiance Fields (NeRF) (Mildenhall et al., 2021) in the latent space. (Lin et al., 2023; Chen et al., 2023a) optimize efficient mesh representations (Munkberg et al., 2022; Shen et al., 2021) for higher quality generation. (Seo et al., 2023) integrates 3D awareness into 2D diffusion for improving text-to-3D generation consistency. Subsequent studies (Chen et al., 2023b; Lugmayr et al., 2022; Richardson et al., 2023) focus on texturing pre-existing meshes, balancing speed and quality. Despite their innovations, SDS-based methods faced challenges with over-saturation. (Tsalicoglou et al., 2023) proposes a novel method to refine mesh textures for more realistic generations. ProlificDreamer (Wang et al., 2023b) introduces the Variational Score Distillation (VSD) method to mitigate over-saturation effectively. Furthermore, several studies (Shi et al., 2023b; Ye et al., 2023; Zhao et al., 2023; Liu et al., 2023a) explored 3D generation using multi-view diffusion. Concurrently, some research (Huang et al., 2023; Liu et al., 2023c, b; Long et al., 2023; Qian et al., 2023; Raj et al., 2023; Shi et al., 2023a; Tang et al., 2023; Xu et al., 2023; Wu et al., 2023; Melas-Kyriazi et al., 2023) concentrates on reconstructing 3D content from a single image through distillation, achieving high-fidelity textured meshes from 2D diffusion priors. Additionally, some studies (Yu et al., 2023; Erkoç et al., 2023; Nash et al., 2020; Siddiqui et al., 2023) delve into shape generation. (Yu et al., 2023) generates high-quality 3D shapes with Unsigned Distance Field (UDF) through Diffusion models. (Nash et al., 2020; Siddiqui et al., 2023) employ autoregressive models for mesh structure generation. Although some general object generation methods (Qiu et al., 2023; Mildenhall et al., 2021; Poole et al., 2022; Wang et al., 2023b, a; Yu et al., 2023) can produce garments, their practicality in CG environments is limited. Very Recent work, Garment3DGen (Sarafianos et al., 2024) enables users to generate textured 3D garments from single images or text prompts based on a 3D base mesh. These 3D outputs, mostly mesh-based or derived from implicit fields, lack adaptability for fitting different bodies and layering multiple garments, common needs in garment design. Furthermore, the textures, typically produced via optimization or multi-view reconstruction, are often low-resolution and blurry, neglecting the structured UV mapping of garments, resulting in poor topology challenging for subsequent CG processing.
3. Sewing Pattern Generation
Figure 2. Overview of our SewingGPT pipeline. We quantize sewing patterns to the sequence of tokens and adopt a GPT-based architecture to generate the tokens autoregressively. Our SewingGPT enables users to generate highly diverse and high-quality sewing patterns under text prompt guidance.
Figure 3. Details of our quantization. We present an example of a part of a sleeveless dress, including a skirt panel (Panel 1) and a top panel (Panel 2). Assuming $N_{1}<N_{2}=K$, we require zero-padding for tokens from Panel 1.
@@ 3. Sewing Pattern Generation
3.1. Sewing Pattern Quantization
@@
We utilize the sewing pattern templates from (Korosteleva and Lee, 2022), which cover a wide variety of garment shapes. Each sewing pattern includes $N_{P}$ panels $\{P_{i}\}_{i=1}^{N_{P}}$ and stitching information $S$. Each panel $P_{i}$ forms a closed 2D polygon with $N_{i}$ edges $\{E_{i,j}\}_{j=1}^{N_{i}}$. Each edge $E_{i,j}$ consists of four parameters $(v_{x},v_{y},c_{x},c_{y})$, where $(v_{x},v_{y})$ represents the edge’s start point, and $(c_{x},c_{y})$ represents the control point of the Bezier curve. Since the panels form closed polygons, we do not need to store the edges’ endpoints. The 3D placement of each panel is indicated by rotation quaternion $R_{i}\in\text{SO}(3)$ and translation vector $T_{i}\in\mathbb{R}^{3}$. For stitching information, we utilize per-edge stitch tags $\{S_{i,j}\}_{j=1}^{N_{i}}$ and stitch flags $\{U_{i,j}\}_{j=1}^{N_{i}}$ for each panel $P_{i}$, obtained from the stitching information $S$. Each stitch tag $S_{i,j}\in\mathbb{R}^{3}$ is based on the 3D placement of the corresponding edge, and each stitch flag is a binary flag with $U_{i,j}=\{0,1\}$ indicating whether there is a stitch on this edge. We follow a similar approach to that used in (Korosteleva and Lee, 2022), which utilizes Euclidean distance between tags as a similarity measure. To restore the stitching information from stitch tags and stitch flags, we filter out free and connected edges with stitch flags and then compare the stitch tags of all pairs of connected edges.
@@
For each panel, we first utilize a similar data preprocessing approach to that used in (Korosteleva and Lee, 2022), which standardizes all edge vectors and control points to maintain the data within a standard normal distribution, and normalizes its 3D placement to ensure all values are between 0 and 1. Then, we quantize all parameters, subsequently converting them into tokens. Specifically, for panel $P_{i}$, we model all parameters as discrete variables by multiplying predefined constants $C_{E},C_{R},C_{T},C_{S}$ by edge vectors, rotation, translation, and stitching feature vectors respectively, and maintain stitching flags as 0 or 1. We flatten and concatenate $N_{i}$ edges, one rotation quaternion, one translation vector, $N_{i}$ stitching vectors, and $N_{i}$ stitching flags, into a sequence of tokens. We carefully select these constants to offer a good trade-off between maintaining the fidelity of sewing patterns and managing the vocabulary size. Overall, we can represent the quantization process as
(1) $\mathcal{T}(P_{i},S)=C_{E}\{E_{i,j}\}_{j=1}^{N_{i}}\oplus C_{R}R_{i}\oplus C_{T}T_{i}\oplus C_{S}\{S_{i,j}\}_{j=1}^{N_{i}}\oplus\{U_{i,j}\}_{j=1}^{N_{i}}$
where we denote $\mathcal{T}$ as the quantization function, and $\oplus$ as the linear concatenation of tokens. These tokens are then formed into a linear sequence. We set a maximum limit, denoted as $K$, for the number of edges in each panel. To maintain a uniform token count across panels, we apply zero-padding to panels with $N_{i}<K$. Subsequently, all panels are flattened and merged into a single sequence, starting with a start token and ending with an end token. Owing to the uniformity of token counts for each panel, inserting padding tokens between panels is not required. Consequently, the resultant sequence, as illustrated in Figure 3, denoted as $\mathcal{F^{\text{seq}}}$, spans a length of $L_{t}=(8K+7)N_{P}$, with each token denoted by $f_{n}$ for $n=1,\ldots,L_{t}$. Eventually, we can represent it as
(2) $\mathcal{F^{\text{seq}}}=\{\mathcal{T}(P_{i},S)+C\}_{i=1}^{N_{P}},$
where C is a constant to ensure all tokens are non-negative.
3.2. Generation with Autoregressive Model
Utilizing GPT-based architectures, we adopt a decoder-only transformer to generate token sequences for sewing patterns. Inspired by PolyGen (Nash et al., 2020), we design the triple embedding for each input token: positional embedding, denoting which panel it belongs to; parameter embedding, classifying the token as edge coordinates, rotation, translation, or stitching feature vectors; and value embedding for the quantized sewing pattern values. The source tokens are then input into the transformer decoder to predict the probability distribution of the next potential token at each step. Consequently, the objective is to maximize the log-likelihood of the training sequences:
(3) $\mathcal{L}=-\prod_{i=1}^{L_{t}}p\left(f_{i}\mid f_{<i};\theta\right).$
@@ 3.2. Generation with Autoregressive Model
To guide the sewing pattern generation, our model integrates cross-attention with text-conditioned embeddings h. Initially, we utilize the CLIP model to obtain the CLIP embedding from input text prompts. Then, we project it into a feature embedding through a trainable compact Multilayer Perceptron (MLP) to condense the dimensionality of CLIP embeddings, matching the Transformer’s dimensionality. This approach also boosts memory efficiency and inference speed. Subsequently, the Transformer decoder conducts cross-attention with the feature embedding (Li et al., 2022). We train the model with pairwise data, facilitating condition-specific token generation.
Figure 4. Examples of our data captions. We utilize the rendered images and ask GPT-4V with the designed prompt for detailed captions.
Figure 5. Overview of our entire DressCode pipeline for customized garment generation. We employ a large language model to obtain shape prompts and texture prompts with natural language interaction and utilize the SewingGPT and a fine-tuned Stable Diffusion for high-quality and CG-friendly garment generation.
3.3. Implementation Details